For math I have to:
Creat an equation for a line perpendicular to y=-2/3x+4
So how the heck do I do this????
I feel like I should know!?
Gracias, Sofia 😃😃😃😃😃😃

Answer by stanbon(75887)

Creat an equation for a line perpendicular to y=-2/3x+4
------
For 2 lines to be perpendicular the product of their slopes must be -1.
-----------------
The given line has slope = -2/3
---
The new equation must have slope = 3/2
-------
Example of such an equation is::
y = (3/2)x
----------------
Cheers,
Stan H.
-----------------